OpenStudy (anonymous):

The area (A) of a ring is A = π r2 - π s2. Find the area A if π = 3.14, r = 6 and s = 4. A = 62.8 A = 197.192 A = 31.4 A = 6.28

OpenStudy (zarkon):

A

OpenStudy (anonymous):

Don't even need a calculator... pi r^2 - pi s^2 = pi (r^2-s^2) = pi (r+s)(r-s) = 3.14 *10*2 = 62.8